f(x) = x - + 5x -4 b) f(x) = xsenx COS X f (x) 1-senx d) f(x) = xe -5e e) F(x)= In 3

derivated the following exercises
Answer #1

SOLUTION :


a.


f(x) = x^2 - 2/x + 5x - 4


 f’ (x) 

= 2x - 2(-1)/x^2 + 5 

= 2x + 2/x^2 + 5 (ANSWER).


b.


f(x) = x^3 sen x  , I think it is x^3 sin x 


Accordingly,


f’ (x)

= x^3 d/dx (sin x) + d/dx (x^3) sin x

= x^3 cos x + 3x^2 sin x (ANSWER).


c.


f(x) = cos x / (1 - sin x) 


f’ (x) 

= cos x d/dx (1 - sin x)^(-1) + d/dx (cos x) / (1 - sin x)

= - cos x / (1 - sin x)^2 * (- cos x) + (- sin x) / (1 - sin x)

= cos^2 x / (1 - sin x)^2 - sin x / (1 - sin x)

= [ cos^2 x - sin x (1 - sin x) ] / (1 - sin x)^2

= [ cos^2 x - sin x + sin^2 x ] / (1 - sin x)^2

= (1 - sin x) / (1 - sin x)^2

= 1 / (1 - sin x) (ANSWER).


d. 


f(x) = xe^x - 5e^x 


f’ (x)

= xe^x + e^x - 5e^x

= xe^x - 4e^x (ANSWER).


e.


f(x) = ln(5/x) = ln(5) - ln(x)


f’ (x)

= 0 - 1/x 

= - 1/x  (ANSWER).

answered by: Tulsiram Garg
Answer #2

Solutions (al finl= n² - 2 / +50-4 differentiate above with respect tox f(m) S ( 13 +54-9) avr229 com flimi= 30²7 (-2) +-11 5(el fine art senn Com септ flame = df con ) (1-son my focess.com (1-sena (-Cosn) -sknx (resencal)- Carn (I-senx)2 - senxt sin

sena senx ) ? (1-sena ) col frals **- seat flame = f (nel seny find a xent en sen el finis ln() fimis (-5*2) femi=-22